OpenStudy (anonymous):

6- 3/12 i got 23/4 but it's incorrect

OpenStudy (anonymous):

\[6-\frac{ 3 }{ 12 }\]

OpenStudy (anonymous):

i think you literally have a learning disability.

jimthompson5910 (jim_thompson5910):

6- 3/12 6(12/12)- 3/12 72/12 - 3/12 (72 - 3)/12 69/12 23/4 So there has to be a typo

OpenStudy (anonymous):

I re- wrote it @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes my work is based off what you wrote

jimthompson5910 (jim_thompson5910):

try writing it as a mixed number

jimthompson5910 (jim_thompson5910):

or in decimal form

OpenStudy (anonymous):

i also got that tooo

OpenStudy (anonymous):

2

OpenStudy (anonymous):

you divide 12 from 3 and you get 4 and you subtract 6 to 4 and you get 2

OpenStudy (anonymous):

\[6-\frac{ 3 }{ 12 }\]

OpenStudy (anonymous):

nop @ceasy007 its incorrect

jimthompson5910 (jim_thompson5910):

does it tell you type in a mixed number?

OpenStudy (anonymous):

\[6-\frac{ 3 }{ 12 }\] @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes i know the problem

jimthompson5910 (jim_thompson5910):

that's all it says though?

OpenStudy (anonymous):

it says evaluate the expression @jim_thompson5910

jimthompson5910 (jim_thompson5910):

well it's 23/4 which as a mixed number is 5 3/4

jimthompson5910 (jim_thompson5910):

in decimal form it is 5.75

OpenStudy (anonymous):

yes totally forgot about that @jim_thompson5910

jimthompson5910 (jim_thompson5910):

which one worked?

OpenStudy (anonymous):

actually the answer i put in the question 23/4 which you also gave me.

OpenStudy (mathstudent55):

First, reduce the fraction: 6 - 3/12 = 6 - 1/4 = 6/1 - 1/4 Now find a common denominator: the LCD of 1 and 4 is 4 24/4 - 1/4 = 23/4 = 5 3/4

jimthompson5910 (jim_thompson5910):

oh, so why didn't it work the first time?

jimthompson5910 (jim_thompson5910):

how odd

OpenStudy (anonymous):

i thought i would only have to go 23/4 but you can reduce that to 5 3/4

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

it doesn't reduce, it converts to that

OpenStudy (anonymous):

yes that's what i meant

jimthompson5910 (jim_thompson5910):

i gotcha